OpenStudy (anonymous):

the % H in HCI=? a.2.8 b.3.77 c.5 d.97.2

OpenStudy (anonymous):

Mass of one hydrogen atom: 1.01 amu Mass of one chlorine atom: 35.45 amu Mass percent Cl in HCl = 35.45 / (1.01 + 35.45) = 35.45 / 36.46 = 97.2%

OpenStudy (anonymous):

So to get %H subtract %CL from 100 so what is 100-97.2ish

OpenStudy (anonymous):

2.8 thanks

OpenStudy (anonymous):

Yep.